HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

VUV-FT SPECTROSCOPY OF THE a^{3}Π − X^{1}Σ^{+} CAMERON BAND SYSTEM IN THE CO MOLECULE

A. STASIK, Materials Spectroscopy Laboratory, Institute of Physics, University of

Rzeszów, Poland; A. N. HEAYS, J. Heyrovsky Institute of Physical Chemistry, Prague, Czech Republic; W. UBACHS, Department of Physics and Astronomy, and

LaserLaB, Vrije Universiteit, Netherlands; R. W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, USA; M. I. MALICKA, The Faculty of Mathematics and Applied Physics, Rzeszów University of Technology,

Rzeszów, Poland; N. DE OLIVEIRA, Synchrotron SOLEIL, France; S. RYZNER,

W. SZAJNA, \underline{R. HAKALLA} , Materials Spectroscopy Laboratory, Institute of Physics, University of Rzeszów, Poland

The study of carbon monoxide is of considerable importance in modelling the evolution of nebulae, stars, planets and exoplanets, as well as entire galaxies, due to the second-largest abundance of this molecule in the universe after the molecular hydrogen.^{1} Cameron system, a^{3}Π − X^{1}Σ^{+} is the subject of very intense study to understand of a daylight of the Mars atmosphere.^{2}^{,}^{3}^{,}^{4}

This work is aimed to obtain the highest energy, non-perturbed band among all the hot bands of the a^{3}Π − X^{1}Σ^{+} Cameron system in the CO molecule, namely the (3, 0), as well as to precisely analyze the a^{3}Π(v = 3) vibrational level. For this purpose, the VUV-FT absorption technique was used by employing the wave-front-division spectrometer working as the end station on the DESIRS beamline (SOLEIL synchrotron).^{5}^{,}^{6} The wavenumbers of the 9 rotational branches of the band under consideration were thus obtained. Based on the effective Hamiltonian and using the PGOPHER program^{7}, the accurate rotational and Λ-doubling molecular constants as well as spin-orbit, spin-rotation, and off-diagonal spin-spin coupling constants of the a^{3}Π(v = 3) level were obtained. The ro-vibronic terms of a^{3}Π(v = 3), based on the experimental transition frequencies, were calculated.

1doi:10.1146/annurev.astro.42.053102.134010, E. F. Van Dieshoeck, Annual Review of Astronomy and Astrophysics, 42, 119-167, (2004).

2doi:10.1016/j.pss.2011.08.010, S.K. Jain et al., Planetary and Space Science, 63-64, 110-112 (2012).

3doi:10.1016/j.icarus.2014.09.051, A. Stiepen et al., Icarus, 245, 295-305, (2015).

4doi:10.1016/j.icarus.2015.09.023, L. Soret et al., Icarus, 264, 398-406, (2016).

5doi:10.1103/PhysRevA.84.052509, A. J. de Nijs et al., Physical Review A, 84, 052509, (2011).

6doi:10.1107/S1600577516006135, N. de Oliveira et al., Journal of Synchrotron Radiation, 23, 887-900, (2016).

7doi:10.1016/j.jqsrt.2016.04.010, C. Western, Journal of Quantitative Spectroscopy and Radiative Transfer, 186, 221-242, (2017).

p-number: p076  Submitted on Mon, 12 Jun 2023 11:09:00 +0200